# Impact of data usage for forecasting on performance of model predictive control in buildings with smart energy storage

### v1

@@ Abstract
Data is required to develop forecasting models for use in Model Predictive Control (MPC) schemes in building energy systems. However, data usage incurs costs from both its collection and exploitation. Determining cost optimal data usage requires understanding of the forecast accuracy and resulting MPC operational performance it enables. This study investigates the performance of both simple and state-of-the-art machine learning prediction models for MPC in a multi-building energy system simulation using historic building energy data. The impact of data usage on forecast accuracy is quantified for the following data efficiency measures: reuse of prediction models, reduction of training data volumes, reduction of model data features, and online model training. A simple linear multi-layer perceptron model is shown to provide equivalent forecast accuracy to state-of-the-art models, with greater data efficiency and generalisability. The use of more than 2 years of training data for load prediction models provided no significant improvement in forecast accuracy. Forecast accuracy and data efficiency were improved simultaneously by using change-point analysis to screen training data. Reused models and those trained with 3 months of data had on average 10% higher error than baseline, indicating that deploying MPC systems without prior data collection may be economic.
@@
Impact of data usage on forecast accuracy and MPC operational performance quantified
@@ 1.1 Background
The operation of building systems accounts for 30% of final energy consumption and 26% of energy-related carbon emissions globally iea2023TrackingCleanEnergy . Thus, decarbonizing building energy usage is necessary to achieve the targets of net-zero carbon emissions by 2050 ccc2020SixthCarbonBudget . The use of distributed generation and storage technologies in building energy systems can enable substantial reductions in operational emissions aminitoosi2022BuildingDecarbonizationAssessing ; oshaughnessey2021DemandSideOpportunityRoles ; zhu2023ReviewDistributedEnergy . Additionally, smart energy storage systems can reduce the impact of building energy usage on the electrical grid by allowing energy flexibility through demand-side management and demand response vazquez-canteli2020MARLISAMultiAgentReinforcement , which is of particular importance in light of the expected electrification of heating demands leibowicz2018OptimalDecarbonizationPathwaysa . The effectiveness of these systems, quantified by operational performance metrics of the building energy system such as total electricity cost, incurred carbon emissions, and measures of grid impact, is determined by the ability of the storage control scheme to arbitrage energy and alter the timing of net energy usage to meet these operational goals. Hence, the development of performant storage control strategies has received substantial research attention in recent years drgona2020AllYouNeed ; wang2020ReinforcementLearningBuilding ; kathirgamanathan2021DatadrivenPredictiveControl .
Model Predictive Control (MPC) is a prominent control methodology in the building energy systems literature. Its application to smart energy storage systems in buildings has been widely studied zhou2021IncorporatingDeepLearning ; deng2023EvaluationDeployingDatadriven ; souto2018ScenarioBasedDecentralizedMPC ; touretzky2014IntegratingSchedulingControl ; lee2020ModelPredictiveControl , and has been found to provide substantial operational performance improvements compared to Rule-Based control (RBC) lee2020ModelPredictiveControl ; erfani2021AnalysisImpactPredictive ; oldewurtel2012UseModelPredictive , the predominant technology used in existing real-world systems. MPC is shown to yield equivalent performance to Reinforcement Learning control (RLC) zhan2023ComparingModelPredictive , the other major control methodology studied in current literature. Review papers wang2022ScienceMappingApproach ; thieblemont2017PredictiveControlStrategies ; farrokhifar2021ModelPredictiveControl provide a thorough overview of the applications of MPC to building energy systems with distributed generation and storage.
@@ 1.2 Forecasting models for MPC
A broad range of time series forecasting methodologies have been studied for the prediction of operational conditions for building energy management sun2020ReviewThestateoftheartDatadriven , however recent literature has focused on the use of machine learning based methods zhang2021ReviewMachineLearning ; chou2018ForecastingEnergyConsumption ; dai2023ComparisonDifferentDeep ; rahman2018PredictingElectricityConsumption ; fan2019AssessmentDeepRecurrent due to their promising performance. Successes in other fields have prompted the investigation of large-scale, high-complexity model architectures choi2023PerformanceEvaluationDeep ; dai2023CityTFTTemporalFusion . However, doubt remains as to whether such complex models are appropriate for use in the context of MPC for building energy systems zeng2023AreTransformersEffective ; bunning2022PhysicsinformedLinearRegression ; bunning2021ComparingMachineLearning , due to computational and data availability limitations in practical systems. As such, this work considers both simple and high-complexity, state-of-the-art machine learning forecasting models.
1.3 Study of impact of data usage on forecast accuracy and MPC performance
@@
As machine learning methods are purely data-driven, “black box” prediction models, achieving accurate forecasts requires the availability of training data that is representative of the building energy system for which the model will be used. However, acquiring representative training data is both challenging and costly. Whilst the importance of data availability and the impact of data usage on forecast accuracy are widely acknowledged in the literature kathirgamanathan2021DatadrivenPredictiveControl ; lee2020ModelPredictiveControl ; wang2022ScienceMappingApproach ; choi2023PerformanceEvaluationDeep ; zhan2021DataRequirementsPerformance , few works study the role of data in enabling good operational performance for MPC in building energy systems.
@@
Determining cost optimal data collection strategies to support the development of forecasting models for MPC in buildings requires understanding of the trade-off between data usage, and its associated costs, and the operational performance achieved by the controller. This impact of data usage for forecasting on MPC operational performance can be considered two stages, by firstly studying the relationship between data usage and forecast accuracy of the resulting models, and then the sensitivity of the controller operational performance to the accuracy of forecasts.
@@
Only a single previous study investigating the impact of data usage on the accuracy of forecasting models for building energy management could be identified. This work, choi2023PerformanceEvaluationDeep , compares the prediction accuracy of deep learning architectures for forecasting thermal loads and building zone temperatures over varying training dataset sizes. It finds that increasing the training data volume does not always improve prediction accuracy due to the strong seasonality of building energy behaviours, but that the usage of additional data with good similarity to the test dataset greatly improves forecast accuracy. A limitation of this study is that it analyses only a single year of thermal load and zone temperature data, synthesised from a building energy model and historical weather data, meaning that a limited range of training dataset sizes, from 3 to 9 months, is considered. As a result, training data preceding the test data by one year, which due to seasonality is likely to have the highest similarity and so greatest value, is unavailable, meaning the study of the benefit of additional training data usage is incomplete.
@@
The impact of forecast accuracy on MPC operational performance has been studied to a limited extent. erfani2021AnalysisImpactPredictive compares the use of various classical and machine learning forecasting models in a common MPC framework, quantifying both the forecast accuracy and resulting operational performance of MPC. However, the models studied all achieve comparable forecast accuracies and operational performances, meaning limited insight can be gained into the relationship between the two. Further, as comparison is made between model types, these results cannot be used to assess the benefit of forecast improvement for any individual model. oldewurtel2012UseModelPredictive shows that the use of more accurate, external weather forecasts improves MPC operational performance, but does not quantify the forecast accuracies for comparison. In bartolucci2019HybridRenewableEnergy , the prediction accuracy and corresponding operational performance of MPC for two forecasts of electrical load with different levels of synthetic noise are quantified. Finally, enriquez2016SolarForecastingRequirements provides the most complete study, investigating the variation of operational performance with the noise amplitude of synthetic forecasts of temperature and solar irradiance. However, as the forecast accuracy of the synthetic predictions is not quantified, these results cannot be compared to the performance of practical prediction models and used to assess the benefits of forecast accuracy improvements.
@@
The direct impact of training data usage on operational performance is studied in savadkoohi2023FacilitatingImplementationNeural . It computes the operational performance achieved by a neural network based building thermal controller as the size of the training dataset varies, and finds that negligible performance improvements are obtained when using more than 8 months of training data. Whilst a comprehensive quantification of the relationship between training data volume and operational performance achieved by the specific control scheme studied is provided, the results are specific to the atypical controller architecture used, which does not include the explicit forecasting of operational variables used in typical MPC schemes. Additionally, other aspects of data usage, such as the inclusion of additional data variables, and the reuse of data from existing buildings, are not considered.
@@
Questions of the impact of data usage on MPC performance in buildings, and the optimal data collection strategies to support model development, have been explored in the System Identification field zhan2021DataRequirementsPerformance ; balali2023EnergyModellingControl ; zhang2023InvestigationsMachineLearningbased ; erfani2023LinkingDatasetQuality ; zhan2022ImpactOccupantRelated ; zhan2022ModelcentricDatacentricPractical in the context of developing accurate system dynamics models for MPC, termed ‘control-oriented models’. Works have investigated the data requirements of different modelling approaches zhan2021DataRequirementsPerformance ; balali2023EnergyModellingControl ; zhang2023InvestigationsMachineLearningbased , the impact of data resolution on model accuracy erfani2023LinkingDatasetQuality , the impact of model prediction accuracy on operational performance zhan2022ImpactOccupantRelated , and cost optimal data collection strategies to support model development zhan2022ModelcentricDatacentricPractical .
@@ 1.4 Research objectives & novel contributions
In the existing literature, the impact of data usage on the prediction accuracy of forecasting models for building energy management has been studied to a very limited extent. Additionally, there has been no study of the trade-off between data usage and the operational performance of MPC for battery scheduling in systems with distributed generation and storage, understanding of which is necessary to properly prioritise expenditure on data collection for smart energy storage systems. This study aims to address this research gap by quantifying the impacts of data usage on both the prediction accuracy and operational performance of MPC using simple and high-complexity, state-of-the-art machine learning based prediction models. Simulation of a multi-building energy system with distributed solar generation and battery storage using historic building load measurement data is used as a case study. The main objectives of this study are:
@@
    Investigate the trade-off between data usage and forecast accuracy for the following data efficiency measures: reuse of prediction models, reduction of training data volumes, reduction of model data features, and online model training;
@@
    Propose strategies for improving prediction performance when selecting models for reuse and selecting data to exclude when reducing training data volumes; and
@@
The key contribution of this work is the combined study of the impact of data usage on both forecast accuracy and the resulting operational performance of MPC. This is important as it allows energy system designers to assess the trade-off between the cost of data usage for forecasting and the operational benefits it provides. The impact of aspects of data usage on forecast accuracy not yet studied in the context of building energy management are also investigated, specifically the reuse of prediction models, selection of model data features, and online model training. Two strategies for improving the efficacy of data usage for building load forecasting are proposed: a load profile similarity metric for selecting prediction models for reuse, and a change-point detection based methodology for screening training data to improve prediction performance whilst reducing model training time. Further, a long duration (10 year) historic building energy dataset is used to conduct these experiments, allowing performance to be evaluated on a multi-year scenario, providing more robust testing than existing studies.
The remainder of this work is structured as follows. Section 2 describes the simulation framework and data used to perform the experiments, as well as the models tested, and how their performance is evaluated. In Section 3.1, the forecast accuracy of the prediction models trained without data usage limitations is evaluated to provide a baseline, and model performance is compared. Model generalisation for load prediction between buildings is then tested in Section 3.2, to assess whether model reuse is a viable strategy for reducing data collection requirements for new smart energy storage systems. A load profile similarity metric based on the Wasserstein distance between fPCA coefficient distributions is proposed, and its efficacy as a criterion for selecting models for reuse is tested. Section 3.3 studies the impact of various aspects of data usage on forecasting accuracy. The effect of the volume of data used for model training is studied in Section 3.3.1 to support decision making on the quantity of data that should be collected for model development. A change-point detection based methodology for screening training data is proposed, and its ability to improve prediction accuracy whilst reducing data usage is investigated. The selection of model data features and use of online model training are considered in Sections 3.3.2 and 3.3.3 respectively. Section 3.4 contextualises the study of model forecast accuracy in the building energy system control task by quantifying the relationship between forecast accuracy and the resulting MPC operational performance for synthetic noisy forecasts. Finally, conclusions are drawn in Section 4.
@@ 2.1 Smart building control simulation framework
To study the impacts of data usage on forecasting and MPC performance in the context of building energy systems, a simulation environment for an example system was developed. In this test case, MPC is used to schedule battery storage operation in a multi-building energy system with distributed storage and solar generation, to reduce the electricity price, carbon emissions, and grid impact associated with meeting the electrical demand of the buildings. The CityLearn vazquez-canteli2019CityLearnV1OpenAI ; vazquez-canteli2020CityLearnStandardizingResearch building energy control framework is used to simulate the behaviour of the building energy system, and provide the required data to a Linear Program based MPC implementation. A schematic of the energy flows within the simulated multi-building energy system is provided in Fig. 1.
@@
The Linear Program formulation used in the MPC scheme is described by Eq. 2.1, with Table 1 providing descriptions of the parameters. At each time step the optimised control actions, ${E_{i}}^{*}[\tau{=}0]$, are taken. The optimisation objective is comprised of three weighted components, which correspond to the cost of grid electricity consumed by the buildings (assuming no net electricity metering), the embodied carbon emissions associated with the grid electricity, and the ramping of the overall grid electrical demand which represents the grid impact. The three components are normalised by the values they would have if no battery storage were present in the buildings, denoted by $\widetilde{O}^{k}$ for component $k$, lower bounded at 1. This clipping is performed to prevent ill-conditioning of the objective when the no-storage objective values are small.
$\begin{split}\min&\qquad\frac{\gamma^{p}\,\,{\sum}_{\tau}\,\,p^{e}[t{+}\tau]\,{\sum_{i}}\left(\max\left[0,\,E_{i}^{b}[t{+}\tau]\right]\right)}{\raisebox{-5.16663pt}{$\max\left[1,\widetilde{O}^{p}\right]$}}\,\,+\,\,\frac{\gamma^{c}\,\,{\sum}_{\tau}\,\,c[t{+}\tau]\,{\sum_{i}}\left(\max\left[0,\,E_{i}^{b}[t{+}\tau]\right]\right)}{\raisebox{-5.16663pt}{$\max\left[1,\widetilde{O}^{c}\right]$}}\\[4.30554pt] &\qquad\qquad\qquad\qquad+\,\frac{\gamma^{r}\,\,{\sum}_{\tau}\,\,\mathinner{\!\left\lvert\bigg{(}{\sum_{i}}\,E_{i}^{b}[t{+}\tau]\bigg{)}-\bigg{(}{\sum_{i}}\,E_{i}^{b}[t{+}\tau{-}1]\bigg{)}\right\rvert}}{\raisebox{-5.16663pt}{$\max\left[1,\widetilde{O}^{r}\right]$}}\end{split}$ (1)
@@
$E_{i}[t]$ kWh
Energy intake to battery unit in building $i$ at time $t$
$\textrm{SoC}_{i}[t]$ kWh
State-of-charge of battery unit in building $i$ at time $t$
@@
$\Delta t$ hrs
Time step of simulation data
$C^{s}_{i}$ kWh
Energy capacity of battery unit in building $i$
$C^{\textrm{pv}}_{i}$ kWp
Peak power capacity of solar PV unit in building $i$
$\eta_{i}$
Round-trip efficiency of battery unit in building $i$
$P^{\textrm{max}}_{i}$ kW
Power capacity of battery unit in building $i$
$\textrm{SoC}_{i}^{t}$ kWh
State-of-charge of battery unit in building $i$ at time $t$
$L_{i}[t]$ kWh
Electrical demand of building $i$ at time $t$
$g[t]$ kW/kWp
Normalised generation power from solar PV at time $t$
$p^{e}[t]$/kWh
Grid electricity price at time $t$
[…]
Carbon intensity of grid electricity at time $t$
$\gamma^{p}$
Fraction of objective from electricity cost component
$\gamma^{c}$
[…]
$\gamma^{r}$
Fraction of objective from grid ramping component

### v2

# Impact of data for forecasting on performance of model predictive control in buildings with smart energy storage

@@ Abstract
Data is required to develop forecasting models for use in Model Predictive Control (MPC) schemes in building energy systems. However, data is costly to both collect and exploit. Determining cost optimal data usage strategies requires understanding of the forecast accuracy and resulting MPC operational performance it enables. This study investigates the performance of both simple and state-of-the-art machine learning prediction models for MPC in multi-building energy systems using a simulated case study with historic building energy data. The impact on forecast accuracy of measures to improve model data efficiency are quantified, specifically for: reuse of prediction models, reduction of training data duration, reduction of model data features, and online model training. A simple linear multi-layer perceptron model is shown to provide equivalent forecast accuracy to state-of-the-art models, with greater data efficiency and generalisability. The use of more than 2 years of training data for load prediction models provided no significant improvement in forecast accuracy. Forecast accuracy and data efficiency were improved simultaneously by using change-point analysis to screen training data. Reused models and those trained with 3 months of data had on average 10% higher error than baseline, indicating that deploying MPC systems without prior data collection may be economic.
@@
Impact of data on forecast accuracy and MPC operational performance quantified
@@ 1.1 Background
The operation of building systems accounts for 30% of final energy consumption and 26% of energy-related carbon emissions globally [1]. Thus, decarbonizing building energy usage is necessary to achieve the targets of net-zero carbon emissions by 2050 [2]. The use of distributed generation and storage technologies in building energy systems can enable substantial reductions in operational emissions [3, 4, 5]. Additionally, smart energy storage systems can reduce the impact of building energy usage on the electrical grid by allowing energy flexibility through demand-side management and demand response [6], which is of particular importance in light of the expected electrification of heating demands [7]. The effectiveness of these systems, quantified by operational performance metrics of the building energy system such as total electricity cost, incurred carbon emissions, and measures of grid impact, is determined by the ability of the storage control scheme to arbitrage energy and alter the timing of net energy usage to meet these operational goals. Hence, the development of performant storage control strategies has received substantial research attention in recent years [8, 9, 10].
Model Predictive Control (MPC) is a prominent control methodology in the building energy systems literature. Its application to smart energy storage systems in buildings has been widely studied [11, 12, 13, 14, 15], and has been found to provide substantial operational performance improvements compared to Rule-Based control (RBC) [15, 16, 17], the predominant technology used in existing real-world systems. MPC is shown to yield equivalent performance to Reinforcement Learning control (RLC) [18], the other major control methodology studied in current literature. Review papers [19, 20, 21] provide a thorough overview of the applications of MPC to building energy systems with distributed generation and storage.
@@ 1.2 Forecasting models for MPC
A broad range of time series forecasting methodologies have been studied for the prediction of operational conditions for building energy management [22], however recent literature has focused on the use of machine learning based methods [23, 24, 25, 26, 27] due to their promising performance. Successes in other fields have prompted the investigation of large-scale, high-complexity model architectures [28, 29]. However, doubt remains as to whether such complex models are appropriate for use in the context of MPC for building energy systems [30, 31, 32], due to computational and data availability limitations in practical systems. As such, this work considers both simple and high-complexity, state-of-the-art machine learning forecasting models.
1.3 Study of impact of data on forecast accuracy and MPC performance
@@
As machine learning methods are purely data-driven, “black box” prediction models, achieving accurate forecasts requires the availability of training data that is representative of the building energy system for which the model will be used. However, acquiring representative training data is both challenging and costly. The costs of acquiring data include not only the capital costs of installing monitoring systems, but also the costs of digital infrastructure, data processing, system maintenance, and quality assurance required to support data collection. For the case of developing forecasting models for MPC considered in this study, there are additional project costs associated with delaying the installation of the battery system whilst data is gathered. [33] estimates the capital cost of electricity and gas smart-metering for a university campus to be \$0.27/m2, plus an additional \$0.11/m2 for maintenance and supporting IT systems. However, this considers only the cost of collecting data, and neglects the significant costs of training and deploying machine learning models [34], from both the computing and expertise required. Whilst the importance of data availability and the impact of data on forecast accuracy are widely acknowledged in the literature [10, 15, 19, 28, 35], few works study the role of data in enabling good operational performance for MPC in building energy systems.
@@
Determining cost optimal data collection strategies to support the development of forecasting models for MPC in buildings requires understanding of the trade-off between the quantity of data & data features used for model training, and its associated costs, and the operational performance achieved by the controller. This impact of data for forecasting on MPC operational performance can be considered in two stages, by firstly studying the relationship between data and forecast accuracy of the resulting models, and then the sensitivity of the controller operational performance to the accuracy of forecasts.
@@
Only a single previous study investigating the impact of data on the accuracy of forecasting models for building energy management could be identified. This work, [28], compares the prediction accuracy of deep learning architectures for forecasting thermal loads and building zone temperatures over varying training dataset sizes. It finds that increasing the training data length does not always improve prediction accuracy due to the strong seasonality of building energy behaviours, but that the addition of data with good similarity to the test dataset into the training dataset greatly improves forecast accuracy. A limitation of this study is that it analyses only a single year of thermal load and zone temperature data, synthesised from a building energy model and historical weather data, meaning that a limited range of training dataset sizes, from 3 to 9 months, is considered. As a result, training data preceding the test data by one year, which due to seasonality is likely to have the highest similarity and so greatest value, is unavailable, meaning the study of the benefit of additional training data is incomplete.
@@
The impact of forecast accuracy on MPC operational performance has been studied to a limited extent. [16] compares the use of various classical and machine learning forecasting models in a common MPC framework, quantifying both the forecast accuracy and resulting operational performance of MPC. However, the models studied all achieve comparable forecast accuracies and operational performances, meaning limited insight can be gained into the relationship between the two. Further, as comparison is made between model types, these results cannot be used to assess the benefit of forecast improvement for any individual model. [17] shows that the use of more accurate, external weather forecasts improves MPC operational performance, but does not quantify the forecast accuracies for comparison. In [36], the prediction accuracy and corresponding operational performance of MPC for two forecasts of electrical load with different levels of synthetic noise are quantified. Finally, [37] provides the most complete study, investigating the variation of operational performance with the noise amplitude of synthetic forecasts of temperature and solar irradiance. However, as the forecast accuracy of the synthetic predictions is not quantified, these results cannot be compared to the performance of practical prediction models and used to assess the benefits of forecast accuracy improvements.
@@
The direct impact of training data length on operational performance is studied in [38]. It computes the operational performance achieved by a neural network based building thermal controller as the size of the training dataset varies, and finds that negligible performance improvements are obtained when using more than 8 months of training data. Whilst a comprehensive quantification of the relationship between training data volume and operational performance achieved by the specific control scheme studied is provided, the results are specific to the atypical controller architecture used, which does not include the explicit forecasting of operational variables used in typical MPC schemes. Additionally, other aspects of data, such as the inclusion of additional data variables, and the reuse of data from existing buildings, are not considered.
@@
Questions of the impact of data on MPC performance in buildings, and the optimal data collection strategies to support model development, have been explored in the System Identification field [35, 39, 40, 41, 42, 43] in the context of developing accurate system dynamics models for MPC, termed ‘control-oriented models’. Works have investigated the data requirements of different modelling approaches [35, 39, 40], the impact of data resolution on model accuracy [41], the impact of model prediction accuracy on operational performance [42], and cost optimal data collection strategies to support model development [43].
@@ 1.4 Research objectives & novel contributions
In the existing literature, the impact of data on the prediction accuracy of forecasting models for building energy management has been studied to a very limited extent. Additionally, there has been no study of the trade-off between the quantity of data & data features used for model training, and the operational performance of MPC for battery scheduling in systems with distributed generation and storage. Understanding of this trade-off is necessary to properly prioritise expenditure on data collection for smart energy storage systems. This study aims to address this research gap by quantifying the impacts of data on both the prediction accuracy and operational performance of MPC using simple and high-complexity, state-of-the-art machine learning based prediction models. Simulation of a multi-building energy system with distributed solar generation and battery storage using historic building load measurement data is used as a case study. The main objectives of this study are:
@@
    Investigate the trade-off between data and forecast accuracy for the following data efficiency measures: reuse of prediction models, reduction of training data durations, reduction of model data features, and online model training;
@@
    Propose strategies for improving prediction performance when selecting models for reuse and selecting data to exclude when reducing training data durations; and
@@
The key contribution of this work is the combined study of the impact of data on both forecast accuracy and the resulting operational performance of MPC. This is important as it allows energy system designers to assess the trade-off between the cost of data for forecasting and the operational benefits it provides. The impact of aspects of data on forecast accuracy not yet studied in the context of building energy management are also investigated, specifically the reuse of prediction models, selection of model data features, and online model training. Two strategies for improving the efficacy of collected data for building load forecasting are proposed: a load profile similarity metric for selecting prediction models for reuse, and a change-point detection based methodology for screening training data to improve prediction performance whilst reducing model training time. Further, a long duration (10 year) historic building energy dataset is used to conduct these experiments, allowing performance to be evaluated on a multi-year scenario, providing more robust testing than existing studies.
The remainder of this work is structured as follows. Section 2 describes the simulation framework and data used to perform the experiments, as well as the models tested, and how their performance is evaluated. In Section 3.1, the forecast accuracy of the prediction models trained without data duration limitations is evaluated to provide a baseline, and model performance is compared. Model generalisation for load prediction between buildings is then tested in Section 3.2, to assess whether model reuse is a viable strategy for reducing data collection requirements for new smart energy storage systems. A load profile similarity metric based on the Wasserstein distance between functional Principal Component Analysis (fPCA) coefficient distributions is proposed, and its efficacy as a criterion for selecting models for reuse is tested. Section 3.3 studies the impact of various aspects of data on forecasting accuracy. The effect of the volume of data used for model training is studied in Section 3.3.1 to support decision making on the quantity of data that should be collected for model development. A change-point detection based methodology for screening training data is proposed, and its ability to improve prediction accuracy whilst reducing training data durations is investigated. The selection of model data features and use of online model training are considered in Sections 3.3.2 and 3.3.3 respectively. Section 3.4 contextualises the study of model forecast accuracy in the building energy system control task by quantifying the relationship between forecast accuracy and the resulting MPC operational performance for synthetic noisy forecasts. Finally, conclusions are drawn in Section 4.
@@ 2.1 Smart building control simulation framework
To study the impacts of data on forecasting and MPC performance in the context of building energy systems, a case study of an example multi-building building energy system was simulated. In this case study, MPC is used to schedule battery storage operation in the multi-building energy system containing distributed storage and solar generation, as to reduce the electricity price, carbon emissions, and grid impact associated with meeting the electrical demand of the buildings. The CityLearn [44, 45] building energy control framework is used to simulate the behaviour of the building energy system, and provide the required data to a Linear Program based MPC implementation. A schematic of the energy flows within the simulated multi-building energy system is provided in Fig. 1.
The Linear Program formulation used in the MPC scheme is described by Eq. 1, with Table 1 providing descriptions of the parameters. At each time step the optimised control actions, ${E_{i}}^{*}[\tau{=}0]$, are taken. The optimisation objective is comprised of three weighted components, which correspond to the cost of grid electricity consumed by the buildings (assuming no net electricity metering), the embodied carbon emissions associated with the grid electricity, and the ramping of the overall grid electrical demand which represents the grid impact. The three components are normalised by the values they would have if no battery storage were present in the buildings, denoted by $\widetilde{O}^{k}$ for component $k$, lower bounded at 1. This clipping is performed to prevent ill-conditioning of the objective when the no-storage objective values are small.
@@
[…]
@@
$E_{i}[\tau]$ kWh Energy intake to battery unit in building $i$ at time $\tau$ in planning horizon
$\textrm{SoC}_{i}[\tau]$ kWh State-of-charge of battery unit in building $i$ at time $\tau$ in planning horizon
@@
$\Delta t$ hrs Time step of simulation data
$C^{s}_{i}$ kWh Energy capacity of battery unit in building $i$
$C^{\textrm{pv}}_{i}$ kWp Peak power capacity of solar PV unit in building $i$
$\eta_{i}$ Round-trip efficiency of battery unit in building $i$
$P^{\textrm{max}}_{i}$ kW Power capacity of battery unit in building $i$
$\textrm{SoC}_{i}^{t}$ kWh State-of-charge of battery unit in building $i$ at time $t$
$L_{i}[t]$ kWh Electrical demand of building $i$ at time $t$
$g[t]$ kW/kWp Normalised generation power from solar PV at time $t$
$p^{e}[t]$/kWh Grid electricity price at time $t$
$c[t]$ kgCO2/kWh Carbon intensity of grid electricity at time $t$
$\gamma^{p}$ Fraction of objective from electricity cost component
$\gamma^{c}$ Fraction of objective from CO2 emissions component
$\gamma^{r}$ Fraction of objective from grid ramping component